# ORB-SfMLearner: ORB-Guided Self-supervised Visual Odometry with Selective Online Adaptation

### v2

@@ I Introduction
Estimating camera ego-motion from monocular videos is essential for various computer vision and robotics tasks. Especially recently, concentrated research [1, 2] on 3D representation has led to higher demands for camera pose estimation. Traditional methods [3, 4, 5, 6] find matches across frames and restore camera transformations with epipolar geometry, while learning-based self-supervised methods [7, 8, 9] usually infer depth and ego-motion simultaneously and then establish a self-supervised constraint with the photometric reconstruction error. Learning-based methods have been widely studied in recent years due to its fast inference and ability to learn high-level features from data [10, 11].
However, learning-based visual odometry (VO) still faces several challenges that limit its application. First, the accuracy of depth and ego-motion prediction remains inferior to that of traditional methods. Second, due to the black-box nature of neural networks, the decision-making process is not well understood, which reduces trust in the system. Finally, and most importantly, learning-based VO suffers great performance drop on unseen test scene because of the common large domain gap. Even within the domain of autonomous driving, factors such as vehicle speed and weather changes can have a significant impact. We even find that the model’s performance is poor on some already seen training samples that display a certain domain gap, indicating its limited generalization capability and also fitting ability.
@@
Figure 1: The pipeline of the proposed ORB-SfMLearner. The VO DepthNet predicts depth, while the PoseNet predicts the relative pose between two frames after fusing ORB and RGB features through a multi-head cross attention mechanism. The network is trained using the self-supervised reprojection error $\mathcal{L}_{p}$ [7], geometry consistency error $\mathcal{L}_{c}$ [16], and depth map smoothness error $\mathcal{L}_{s}$ [9]. During inference, the network selectively performs online adaptation, learning to utilize the weights most suitable for the current scene, thereby achieving good generalization in challenging conditions, such as foggy weather.
@@
  • We optimize our VO system for generalizability during both training and online adaptation phases. Our online adaptation strategy enables rapid optimization of network parameters based on current data and selects the optimal parameters to output refined prediction.
@@ II Related Works
Visual Odometry is a technique for robots to locate themselves with image stream from visual sensors. As a crucial part in the SLAM system, the principal task for VO is to provide camera pose estimation, also known as the ego motion estimation. In the context of feature-based SLAM [17, 4, 5, 6], previous work such as ORB-SLAM [4] chooses to use ORB features, which offer robust pose recognition ability and efficient real-time tracking. Similar to classical direct VO [17], the core of recent learning-based self-supervised VO is to minimize photometric loss. SfMLearner [7] first proposes to train a PoseNet together with a DepthNet and optimize based on photometric error. Monodepth2 [9] continues this approach but uses ResNet and U-Net architectures for feature extraction and decoding outputs for depth maps and poses. Additionally, Monodepth2 proposes an auto-mask strategy to ignore pixels that remain stationary between frames, which helps reduce the impact caused by synchronized camera and object movement, camera stillness, or weak object textures. SC-SfMLearner [8] further introduces a geometric consistency loss, which calculates reprojection error by transforming the depth maps of adjacent frames based on the estimated pose. This constrains the continuity and structure of the depth maps across frames. These works are key benchmarks for subsequent self-supervised depth and pose estimation research. Building on the core supervision being the photometric error from reprojection, they focus on contributions in areas such as depth map scale consistency and moving object masking and have achieved satisfactory results on the KITTI [18] benchmark.
@@
Explainability in Deep VO has also been explored by some recent work, but mostly focusing on the rationale behind depth prediction. One commonly used method is feature visualization, such as CAM [19] and Grad-CAM [20], to highlight important regions in an image. In addition, Tom et al. [21] treat the DepthNet as a black box and measure the depth prediction giving different input. Specifically, they tested how factors like object position, occlusion, and types affect the model’s predictions. Interestingly, removing the center portion of a car does not significantly impact depth prediction, but if the edges of the car’s bottom are also removed, the DepthNet might fail to recognize the car. For the pose estimation, Sattler et al. [22] revealed that supervised absolute pose regression is essentially more related to image retrieval instead of geometry-based learning. Pose regression methods regress the camera pose of an input image, while self-supervised VO usually takes two consecutive frames and outputs a relative pose estimation. To the best of our knowledge, what the self-supervised PoseNet focuses on remains unexplored.
@@
Generalizability is a critical factor in determining the applicability of learning-based VO in real-world environments. Li et al. [14] use LSTM [23] to extract temporal and spatial information from the input data and enables the network to continuously optimize its parameters based on past experiences. CoVIO [13] continuously updates network weight during inference because of its self-supervised nature, which is similar to our method. They designed a replay buffer, where new frames with cosine similarity to frames in the buffer below a certain threshold are added to the buffer. In contrast, we selectively update parameters based on the self-supervised loss.
@@ III-A Pipeline Overview
We aim to build a VO pipeline that takes advantage of more stable features and adapt itself to overcome variance of testing scenarios. As shown in Fig. 1, with two consecutive frames $(I_{t},I_{t+1})$ from a monocular video as input, our pipeline first augments image data by extracting the ORB features. Then we input the original RGB data and the extracted ORB features into two separate encoders. We apply cross-attention to weight the importance of encoded ORB features relative to RGB features. The fused features are then fed into a decoder to predict the relative 6D camera pose between the two frames. Meanwhile, our DepthNet takes one original image input from $I_{t}$ and $I_{t+1}$ each time, and output the disparity estimation of the current frame.
@@
Finally, in the test phase, both the PoseNet and the DepthNet optimize their parameters based on the current testing snippet. Our online adaptation algorithm rapidly update and choose the most suitable parameters for estimating camera ego-motion and depth. Unlike during training where we avoid overfitting, we actually aim for the model to overfit on the current tested snippet.
@@ III-B Self-supervision Principle for VO
The essence of self-supervised VO lies in reconstructing adjacent frames using the depth and inter-frame relative pose output by the neural network. The similarity between real reconstruction target and the synthesized one reflects the quality of the estimated depth and pose. Given two frames $I_{a}$ and $I_{b}$, the depths $D_{a}$ and $D_{b}$ of them are predicted by the DepthNet and their relative transformation $T_{ab}$ by the PoseNet. We synthesize the reconstructed $I_{b}^{\prime}$ with the differentiable warping process proposed by Zhou et al. [7] and choose $\mathcal{L}_{1}$ and structural similarity (SSIM) loss to construct the photometric loss function $\mathcal{L}_{p}$:
@@ III-D Explain the ORB Guidance
$\text{Output}_{b,n,i,:}=\sum_{j}\text{Attn\_Weights}_{b,n,i,j}\cdot V_{b,n,j,:}.$ (4)
@@
Figure 3: By visualizing the attention weights, a clear pattern emerges: during left or right turns, the regions with high weights also shift accordingly, often pointing toward the end of the road. The two columns on the left are selected from KITTI Odometry sequence 09, and the rightmost column is selected from sequence 07.
@@ III-E Selective Online Adaptation
After pre-training the PoseNet and DepthNet on the source domain, we continue to adapt the model during inference with our proposed selective online adaptation (SOA), as illustrated in Algorithm 1. We first prepare the incoming frames with the proposed ORB-Augmentation. Then we conduct inference with current PoseNet and DepthNet and use the estimated relative pose $P$ and depth $D_{i},D_{i+1}$ to perform differentiable warping (Section III-B). Here, the loss from this self-supervised process is used to update the model parameters. However, we don’t directly iterate to optimize and output the results. Instead, we infer with the updated model to compute the loss again, and only if the loss decreases do we select the updated parameters. We call this strategy selective adaptation. For simplicity, Algorithm 1 only shows the case where each input snippet consists of 2 frames. We can also input more frames and perform differentiable warping on each pair of frames, updating the model online using the average loss.
@@
3:     back up $current\_params$
4:     Pre-process with ORB-Augmentation
5:     for $n=0$ to $k$ do
6:         $D_{i}$ and $D_{i+1}$ from $DepthNet$, $P$ from $PoseNet$($I_{i},I_{i+1}$)
7:         $error_{photometric}\leftarrow D_{i},D_{i+1},I_{i},I_{i+1},P,K$
8:         $error_{geometric}\leftarrow D_{i},D_{i+1},P,K$
9:         $current\_error=error_{photometric}+\alpha\times error_{geometric}$, $\alpha=0.5$
10:         if $current\_error<lowest\_error$ or $n=0$ then
11:              $lowest\_error\leftarrow current\_error$
12:              $best\_params\leftarrow current\_params$
13:         end if
14:         Update $current\_params$ with back propagation
15:     end for
16:     Update DepthNet and PoseNet with $best\_params$, output results inferred with updated models
17:end for
@@ IV-C Camera Ego-motion Estimation
For online adaptation, we input three frames at a time and perform $k=2$ rounds of iteration on these frames. To further validate the effectiveness of our SOA, we conduct an ablation study on this. We test the results on sequences 07 and 09 under the following conditions: (1) varying numbers of iterations, (2) without the selective strategy (optimize the parameters without selecting the best ones based on self-supervised losses), and (3) different numbers of frames input at each step. TABLE II shows the results of online adaptation at 4 different configurations. We observe that the proposed selective adaptation strategy effectively improves the pose estimation precision. In addition, increasing the number of iterations or the snippet sequence length does not necessarily improve prediction accuracy. Therefore, in our other experiments, we adopt the more efficient configuration of using two iterations with a snippet length of three.
@@ IV-D Attention Weights Visualization
In addition to examining the patterns of attention weights between every two consecutive frames in a sequence, we also investigate the impact of the interval between frames on the attention weights. Fig. 5 shows how the attention weights change when setting the $n_{th}$ and $k_{th}$ frames as inputs. We find that when the interval between two frames is small, such as during complete stillness, the highlighted regions in the attention weights are often distributed across the entire image. As the interval increases, the attention focuses more on the end of the road. This pattern aligns with the intuition from Fig. 3 that it is difficult to determine if a car is turning left or right when stationary. In addition, as the frame interval increases too much, the highlighted regions also shift.
@@ IV-E KITTI to vKITTI
We select three weather conditions (i.e., $Clone$, $Fog$, and $Morning$) from vKITTI Scene 01 and 20. The clone condition replicates the weather conditions of KITTI. In addition, we identify real sequences from KITTI for comparison. The pose estimation results are shown in TABLE III. The models perform better on KITTI because its data distribution is similar to that of our training set. However, even the conditions in vKITTI $Clone$ can have significant impact on the accuracy of pose estimation. Our proposed SOA effectively overcomes this domain gap, performing equally well under three distinct weather conditions.
@@ V Conclusion
This paper proposes an ORB-augmented VO with SOA. ORB features guide the network’s attention to key regions, using more stable features from the raw data to predict ego motion. Consequently, our model demonstrates superior pose accuracy across all trajectories in the KITTI odometry dataset. Further, cross-attention module illustrates how ORB features guide the extraction of RGB features, providing a level of interpretability. We observe that regions with higher attention weights correspond to the vehicle’s turning direction. Moreover, our ORB-SfMLearner works well under different domains due to optimal parameters adaptation. Overall, the integration of ORB augmentation has improved the accuracy and explainability of our model’s pose estimation, while the SOA has further enhanced its generalization capabilities.

### v4

@@ I Introduction
Estimating camera ego-motion from monocular videos is essential for various computer vision and robotics tasks. Recent advances in 3D representation learning [1, 2] have also heightened the demand for camera pose estimation. Traditional methods [3, 4, 5, 6] find matches across frames and restore camera transformations with epipolar geometry, while learning-based self-supervised methods [7, 8, 9] usually infer depth and ego-motion simultaneously and then establish a self-supervised constraint with the photometric reconstruction error. Learning-based methods have been widely studied in recent years due to its fast inference and ability to learn high-level features from data [10, 11].
However, learning-based visual odometry (VO) still faces several challenges. First, the accuracy of depth and ego-motion estimation remains inferior to that of traditional methods. Second, due to the black-box nature of neural networks, the decision-making process is not well understood, which reduces trust in the system. Finally, and most importantly, learning-based VO suffers great performance drop on unseen test scene because of the common large domain gap. Even within the domain of autonomous driving, factors such as vehicle speed and weather changes can have a significant impact. We even find that the model’s performance is poor on some already seen training samples that display a certain domain gap, indicating its limited generalization capability and also fitting ability.
@@
Figure 1: The pipeline of the proposed ORB-SfMLearner. The VO DepthNet estimates depth, while the PoseNet estimates the relative pose between two frames after fusing ORB and RGB features through a multi-head cross attention mechanism. The network is trained using the self-supervised reprojection error $\mathcal{L}_{p}$ [7], geometry consistency error $\mathcal{L}_{c}$ [16], and depth map smoothness error $\mathcal{L}_{s}$ [9]. During inference, the network selectively performs online adaptation, learning to use the weights most suitable for the current scene, thereby achieving good generalization in challenging conditions, such as foggy weather.
@@
  • We optimize our VO system for generalizability during both training and online adaptation phases. Our online adaptation strategy enables rapid optimization of network parameters based on current data and selects the optimal parameters to output refined estimation.
@@ II Related Works
Visual Odometry is a technique for robots to locate themselves with image stream from visual sensors. As a crucial part in the SLAM system, the principal task for VO is to provide camera pose estimation, also known as the ego motion estimation. In the context of feature-based SLAM [17, 4, 5, 6], previous work such as ORB-SLAM [4] chooses to use ORB features, which offer robust pose recognition ability and efficient real-time tracking. Similar to classical direct VO [17], the core of recent learning-based self-supervised VO is to minimize photometric loss. SfMLearner [7] was the first to jointly train a PoseNet and a DepthNet and optimize them based on photometric error. Monodepth2 [9] continues this approach but uses ResNet and U-Net architectures for feature extraction and decoding outputs for depth maps and poses. Additionally, Monodepth2 proposes an auto-mask strategy to ignore pixels that remain stationary between frames, which helps reduce the impact caused by synchronized camera and object movement, camera stillness, or weak object textures. SC-SfMLearner [8] further introduces a geometric consistency loss, which calculates reprojection error by transforming the depth maps of adjacent frames based on the estimated pose. This constrains the continuity and structure of the depth maps across frames. These works are key baselines for subsequent self-supervised depth and pose estimation research. Building on the core supervision being the photometric error from reprojection, they contribute to areas such as depth map scale consistency and moving object masking, achieving satisfactory results on the KITTI [18] benchmark.
@@
Explainability in Deep VO has also been explored by some recent works, but mostly focusing on the rationale behind depth estimation. One commonly used method is feature visualization, such as CAM [19] and Grad-CAM [20], to highlight important regions in an image. In addition, Tom et al. [21] treat DepthNet as a black box and evaluate depth output in response to different input variations. Specifically, they tested how factors like object position, occlusion, and types affect the model’s predictions. Interestingly, removing the center portion of a car does not significantly impact the results, but if the edges of the car’s bottom are also removed, the DepthNet might fail to recognize the car. For the pose estimation, Sattler et al. [22] revealed that supervised absolute pose regression is essentially more related to image retrieval instead of geometry-based learning. Pose regression methods regress the camera pose of an input image, while self-supervised VO usually takes two consecutive frames and outputs a relative pose estimation. To the best of our knowledge, what the self-supervised PoseNet focuses on remains unexplored.
@@
Generalizability is a crucial for making learning-based VO applicable in real-world environments. Li et al. [14] use LSTM [23] to extract temporal and spatial information from the input data and enables the network to continuously optimize its parameters based on past experiences. CoVIO [13] continuously updates network weights during inference because of its self-supervised nature, which is similar to our method. They designed a replay buffer, where new frames with cosine similarity to frames in the buffer below a certain threshold are added to the buffer. In contrast, we selectively update parameters based on the self-supervised loss.
@@ III-A Pipeline Overview
We aim to build a VO pipeline that takes advantage of more stable features and adapt itself to overcome variance of testing scenarios. As shown in Fig. 1, with two consecutive frames $(I_{t},I_{t+1})$ from a monocular video as input, our pipeline first augments image data by extracting the ORB features. Then we input the original RGB data and the extracted ORB features into two separate encoders. We apply cross-attention to weight the importance of encoded ORB features relative to RGB features. The fused features are then fed into a decoder to predict the relative 6D camera pose between the two frames. Meanwhile, our DepthNet takes one original image input from $I_{t}$ and $I_{t+1}$ each time, and outputs the disparity estimation of the current frame.
@@
Finally, in the test phase, both the PoseNet and the DepthNet optimize their parameters based on the given input. Our online adaptation algorithm rapidly updates and chooses the most suitable parameters for estimating camera ego-motion and depth. Unlike during training, where we avoid overfitting, we actually aim for the model to overfit on the current tested snippet.
@@ III-B Self-supervision Principle for VO
The essence of self-supervised VO lies in reconstructing adjacent frames using the depth and inter-frame relative pose output by the neural network. The similarity between the real reconstruction target and the synthesized one reflects the quality of the estimated depth and pose. Given two frames $I_{a}$ and $I_{b}$, the depths $D_{a}$ and $D_{b}$ of them are predicted by DepthNet and their relative transformation $T_{ab}$ by the PoseNet. We synthesize the reconstructed $I_{b}^{\prime}$ with the differentiable warping process proposed by Zhou et al. [7] and choose $\mathcal{L}_{1}$ and structural similarity (SSIM) loss to construct the photometric loss function $\mathcal{L}_{p}$:
@@ III-D Explain the ORB Guidance
$\text{Output}_{b,n,i,d}=\sum_{j}\text{Attn\_Weights}_{b,n,i,j}\cdot V_{b,n,j,d}.$ (4)
@@
Figure 3: By visualizing the attention weights, a clear pattern emerges: during left or right turns, the regions with high weights also shift accordingly, often pointing towards the end of the road. The two columns on the left are selected from KITTI Odometry sequence 09, and the rightmost column is selected from sequence 07.
@@ III-E Selective Online Adaptation
After pre-training the PoseNet and DepthNet on the source domain, we continue to adapt the model during inference with our selective online adaptation (SOA), as illustrated in Algorithm 1. We first prepare the incoming frames with the proposed ORB-Augmentation. Then we conduct inference with current PoseNet and DepthNet and use the estimated relative pose $P$ and depth $D_{i},D_{i+1}$ to perform differentiable warping (Section III-B). Here, the loss from this self-supervised process is used to update the model parameters. However, we don’t directly iterate to optimize and output the results. Instead, we infer with the updated model to compute the loss again, and only if the loss decreases do we select the updated parameters. We call this strategy selective adaptation. For simplicity, Algorithm 1 only shows the case where each input snippet consists of 2 frames. We can also input more frames and perform differentiable warping on each pair of frames, updating the model online using the average loss.
@@
3:     Pre-process with ORB-Augmentation
4:     for $n=0$ to $k$ do
5:         Obtain $D_{i}$ and $D_{i+1}$ from $DepthNet$, and $P$ from $PoseNet$($I_{i},I_{i+1}$)
6:         $error_{photometric}\leftarrow D_{i},D_{i+1},I_{i},I_{i+1},P,K$
7:         $error_{geometric}\leftarrow D_{i},D_{i+1},P,K$
8:         $current\_error=error_{photometric}+\alpha\times error_{geometric}$, $\alpha=0.5$
9:         if $current\_error<lowest\_error$ or $n=0$ then
10:              $lowest\_error\leftarrow current\_error$
11:              $best\_params\leftarrow current\_params$
12:         end if
13:         Update $current\_params$ with back propagation
14:     end for
15:     Update DepthNet and PoseNet with $best\_params$, output results inferred with updated models
16:end for
@@ IV-C Camera Ego-motion Estimation
For online adaptation, we input three frames at a time and perform $k=2$ rounds of iteration on these frames. To further validate the effectiveness of our SOA, we conduct an ablation study. We test the results on sequences 07 and 09 under the following conditions: (1) varying numbers of iterations, (2) without the selective strategy (optimize the parameters without selecting the best ones based on self-supervised losses), and (3) different numbers of frames input at each step. TABLE II shows the results of online adaptation at 4 different configurations. We observe that the proposed selective adaptation strategy effectively improves the pose estimation precision. In addition, increasing the number of iterations or the snippet sequence length does not necessarily improve prediction accuracy. Therefore, in our other experiments, we adopt the more efficient configuration of using two iterations with a snippet length of three.
@@ IV-D Attention Weights Visualization
In addition to examining the patterns of attention weights between every two consecutive frames in a sequence, we also investigate how the extent of the frame interval affects the attention weights. Fig. 5 shows how the attention weights vary when setting the $n_{th}$ and $k_{th}$ frames as inputs. We find that when the interval between two frames is small, such as during complete stillness, the highlighted regions in the attention weights are often distributed across the entire image. As the interval increases, the attention focuses more on the end of the road. This pattern aligns with the intuition from Fig. 3 that it is difficult to determine if a car is turning left or right when stationary. In addition, as the frame interval increases too much, the highlighted regions also shift.
@@ IV-E KITTI to vKITTI
We select three weather conditions ($Clone$, $Fog$, and $Morning$) from vKITTI Scene 01 and 20. The clone condition replicates the weather of KITTI. We also find out the corresponding real sequences from KITTI for comparison. The pose estimation results are shown in TABLE III. The models perform better on KITTI because its data distribution is similar to that of our training set. However, even the conditions in vKITTI $Clone$ can have a significant impact on the accuracy of pose estimation. Our proposed SOA effectively overcomes this domain gap and performs equally well under three distinct weather conditions.
@@ V Conclusion
This paper proposes an ORB-augmented VO with SOA. ORB features guide the attention of the network to key regions, and more stable information from the raw data are used to estimate ego motion. Consequently, our model demonstrates superior pose accuracy across all trajectories in the KITTI odometry dataset. Further, cross-attention module illustrates how ORB features guide the extraction of RGB features, providing a level of interpretability. We observe that regions with higher attention weights correspond to the vehicle’s turning direction. Moreover, our ORB-SfMLearner works well under different domains due to optimal parameters adaptation. Overall, the integration of ORB augmentation has improved the accuracy and explainability of our model’s pose estimation, while the SOA has further enhanced its generalization capabilities.